OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x − 4) and (x + 4)(x + 4)

OpenStudy (anonymous):

well \[(a+b)(a-b)=a^2-b^2\]

OpenStudy (anonymous):

and \[(a+b)(a+b)=(a+b)^2\]

OpenStudy (anonymous):

and in this case a=x and b=4 so you can go from there